Question1 If Waterway Company issues 10300 shares of $5 par value common stock for $160200, the account Paid-in Capital in Excess of Par will be credited for $51500. O Paid-in Capital in Excess of Par will be credited for $160200 Common Stock will be credited for $51500 Cash will be debited for $108700.
media%2Ffd1%2Ffd1d6505-b93c-4001-b357-4b
media%2F083%2F08347654-5abd-456e-aa4b-0d
Answer #1

1) Journal entry :

Date accounts & explanation debit credit
Cash 160200
Common Stock 51500
Paid in capital in excess of par value 108700

So answer is c) Common Stock will be credited for $51500

2) Price earning ratio = MPS/EPS

EPS = 168000/56000 = 3

Price earning ratio = 24/3 = 8 Times

So answer is a) 8.0 Times

3) Current ratio = 160000/100000 = 1.6

So answer is c) 1.6:1